# SeFFeC: Semantic Facial Feature Control for Fine-grained Face Editing

### v1

@@ 4.1 Implementation Details
Input to our Transformer encoder is the sequence of vectors $w^{+}\in\mathbb{R}^{512\times 18}$ concatenated with the $512$ dimensional semantic face feature embedding extracted by the embedding layer. The encoder consists of four Transformer encoder layers with four self-attention heads [52] each. The output of the Transformer is input to a linear layer that outputs the semantic manipulation vector $s_{e}$. The scaling network takes the current and target face feature values, $m_{j}$ and $m_{j}^{t}$ for a feature $j$ as well as the feature’s embedding $e_{j}$ as input. It consists of three linear layers with 32, 32, and one hidden unit(s), respectively, with a ReLU activation function after the first two linear layers and outputs the scaling value $k$.
To train SeFFeC, we generate data batches by sampling random vectors $z\sim\mathcal{N}(0,1)$ from a standard normal distribution and generate the corresponding vectors $w^{+}$ and images $I$ using the mapping and synthesis network from a pre-trained StyleGAN2 [24] model. For each image we then sample a face feature index $j$ we want to modify from a uniform distribution $j\sim\mathcal{U}(0,23)$ as well as a desired target feature value $m_{j}^{t}$ from a standard normal distribution $m_{j}^{t}\sim\mathcal{N}(0,1)$. Furthermore, we use the pre-trained SPIGA landmark detector [38] to extract landmarks and calculate the current face feature values $m_{j}$ for each generated image $I$. Since we sample $m_{j}^{t}$ from $\mathcal{N}(0,1)$ we normalise $m_{j}$ to follow a standard normal distribution as well:
@@
where $\mu$ and $\sigma$ are the mean and standard deviation for all face features $j$ calculated on the FFHQ dataset, which our StyleGAN2 model was pre-trained on. This ensures that our defined semantic face features are normalised to the same value range when calculating the loss. Using these generated data samples, we can run SeFFeC to predict $w^{+}_{\text{edit}}$, use the StyleGAN2 synthesis network to generate $I_{\text{edit}}$ and the SPIGA landmark detector to extract the new landmarks and calculate the predicted feature values $m_{j}^{p}$. We calculate the loss as defined in Equation 3 and propagate the gradients back through SPIGA and StyleGAN to finally update the parameters of SeFFeC. We optimise SeFFeC for $10^{5}$ steps using the Adam optimiser [26] with a learning rate of $2^{-5}$ and a batch size of 16. Through empirical testing we set the weighting scalars of the loss function to $\lambda_{\text{pix}}=1$, $\lambda_{\text{feat}}=3$, $\lambda_{\text{SFF}}=0.005$, $\lambda_{\text{reg}}=0.1$ and $\lambda_{\text{cor}}=1$.
Figure 3: Example progressive edits performed with SeFFeC. SeFFeC allows to easily and deterministically perform high-quality progressive edits along many different semantic dimensions with explicit control over the desired target feature values. For each demonstration of the progressive edits, we also show the difference between the first and last image, highlighting which parts of the image changed.
Figure 4: Sample face editing results on a real face image of SeFFeC in comparison with the two state-of-the-art methods GANSpace [16] and InterfaceGAN [42]. The original image was inverted into latent space using the e4e framework [51] and subsequently the degree of smiling was edited. Also shown are the difference images between the edited and original images. SeFFeC is the only method that allows for fine-grained and deterministic control of the degree of smiling without distortions.
@@ 4.2 Qualitative Results
Figure 1 shows sample sequential edits performed using SeFFeC. We first generated faces by sampling random latent vectors and decoded them with StyleGAN2 (left). Afterwards, we sequentially edited specific features by changing the current value to the desired target value as indicated by the arrows. We can see that our method can perform multiple fine-grained edits on the desired features while preserving the visual appearance of the rest of the face, suggesting a high degree of disentanglement. In stark contrast to prior work, we are able to calculate the exact value of our face features at any point and set desired values to perform precise and deterministic editing operations. This allows users to anticipate the outcome of an edit before its execution, e.g. setting the feature for mouth openness to a value of -1.1 always results in a closed mouth when normalised to a standard normal distribution. Note that it is possible to define reasonable bounds for each face feature and subsequently normalise the values to an arbitrary range as we did for our intuitive slider user interface (see supplementary material). In this case the smallest value is zero, indicating a closed mouth, and the largest value is one, indicating a fully opened mouth.
@@
Given that our features are on a continuous scale, SeFFeC naturally enables progressive edits as shown in Figure 3: SeFFeC is able to perform the desired changes with fine-grained control over the desired feature values. For each example, the figure also shows the difference between the first and the last image in the edit sequence. We can see that SeFFeC operates highly localised, as significant changes are only visible in the relevant regions. Other changes visible in the difference images are mostly due to high-frequency features, such as the outline of the face. Minor variations along these lines are visible in the difference images but cannot be noticed in the actual images.
@@
Overall, the examples in Figure 1 and Figure 3 show that our method is able to perform localised and precise face editing for a large variety of facial features. While the semantic face features defined in Table 1 only relate a small number of landmarks to each other, we are able to accurately model complex face dynamics. For example, while the definition for mouth openness only relates the y-coordinate of two landmarks from the centre of the mouth, the progressive editing results for this feature in Figure 3 show how the whole mouth opens naturally. We hypothesise that the main reason for the effectiveness of our semantic face features is within the implicit biases and correlations learned by the StyleGAN2 model. To move the two landmarks used for mouth openness apart, SeFFeC learns to open the whole mouth as this is the only possibility given the learned human face distribution of StyleGAN2.
@@
Figure 4 shows editing results on a real face that we inverted into StyleGAN2’s latent space using the e4e encoder [51]. We edited the facial expression twice: once to show less and once to show more smiling. As the results for SeFFeC in the middle row show, it is also possible to manipulate real faces with high precision, which is generally harder as the inverted latent code of the face might not lay within the well-defined regions of StyleGAN2 (see supplementary material for more real face editing results). Since smiling is one of the few attributes that previous methods are also able to control, we compare our results with GANSpace [16] and InterfaceGAN [42]. We can see that both methods fail to accurately reduce the degree of smiling and start to distort other parts of the face, which is most noticeable in the difference images. When increasing the degree of smiling, all three methods are able to accurately edit the face, while SeFFeC still operates more locally. Additionally, SeFFeC provides more fine-grained control over the mouth as it can edit the mouth shape and mouth openness separately, which would also allow us to add a smile without necessarily opening the mouth. Important to note is that SeFFeC did not require any manually labelled training data, while methods like InterfaceGAN rely on large face datasets with smiling attribute annotations.
@@
Model
Determ. Editing
# Attribute Labels $\downarrow$
Attribute Accuracy $\uparrow$
FID $\downarrow$
ID Score $\uparrow$
GANSpace [16]
0
76.1%
36.89
0.818
InterfaceGAN [42]
30.000
83.2% 32.58
0.802
SeFFeC
0
81.8% 27.15 0.833
@@ 4.3 Quantitative Results
We quantitatively evaluate SeFFeC on the smile edit benchmark [30, 29] and compare its performance against two strong baselines, GANSpace [16] and InterfaceGAN [42], using the same StyleGAN2 model for a fair comparison. The task of this benchmark is to convert faces with neutral expressions into smiling faces. The performance is measured with three metrics: attribute accuracy, which measures whether a face is smiling after editing using an attribute classifier pre-trained on the CelebA [31] dataset, Fréchet Inception Distance (FID) [19], calculated between 4000 edited images and the FFHQ dataset to evaluate image quality [24], and Identity Score (ID Score), which measures if the faces’ identity is preserved after editing by calculating the cosine similarity between embeddings extracted from a pre-trained ArcFace model [9]. To make a neutral face smile with SeFFeC we adjust mouth shape and mouth openness as shown in Figure 4. The results of the smile edit benchmark are reported in Table 2. We find that SeFFeC significantly outperforms GANSpace and InterfaceGAN in terms of FID and ID scores. While InterfaceGAN achieves a slightly higher attribute accuracy than SeFFeC, it is important to note that it was trained on the same 30.000 attribute labels as the classifier used to calculate the accuracy metric. In contrast, SeFFeC did not require any attribute labels and did not explicitly learn to manipulate the smiling attribute, while still being competitive with InterfaceGAN and outperforming GANSpace. Furthermore, unlike previous methods, SeFFeC allows for precise, deterministic control over face shape features. This could potentially be one reason for the significant improvement in FID and ID scores, as SeFFeC can automatically adjust the editing intensity based on the measured values of the face features. This allows SeFFeC  for example, to only slightly modify faces that are almost smiling, whereas previous methods would move a fixed amount along a semantic editing direction, which can lead to unnecessary over-editing and distortion.
@@
Model
Pixel Error $\downarrow$
LIPIPS $\downarrow$
Edit Error $\downarrow$
Entanglement $\downarrow$
$\lambda_{\text{pix}}=0$
0.192
0.258
0.664
0.624
$\lambda_{\text{feat}}=0$
0.041
0.045
0.551
0.238
$\lambda_{\text{reg}}=0$
0.043
0.047
0.549
0.267
$\lambda_{\text{cor}}=0$ 0.040 0.041
0.575
0.225
SeFFeC 0.041 0.045 0.529 0.234
SeFFeC 3x
0.065
0.087
0.318
0.340
@@ 4.4 Ablation Experiments
We perform ablation experiments for different versions of SeFFeC and report four different metrics: pixel error, LIPIPS, edit error and entanglement. The pixel error is defined as the mean absolute error (MAE) between the original and the modified image. While we want this to be as small as possible, it should never be zero as this would indicate no change in the face. LIPIPS [63] is a perceptual similarity metric based on neural network features where a smaller value indicates higher similarity. The edit error is defined as the MAE between the predicted feature value $m_{j}^{\text{p}}$ and the target $m_{j}^{\text{t}}$ for the features $m_{j}$ that we actually want to change. Similarly, entanglement measures the MAE between $m_{j}^{\text{p}}$ and $m_{j}^{\text{t}}$ for all features that were not explicitly changed and should stay close to the original value. We report the results in Table 3, which we calculated by generating 10,000 random images and then performed edits on five random semantic features for each face.
We can see that the model is unstable during training without the pixel loss ($\lambda_{\text{pix}}=0$), which is reflected in the lower performance. Training without the feature loss ($\lambda_{\text{feat}}=0$) slightly increases the edit error, indicating that the feature supervision adds some flexibility compared to pure pixel supervision, improving editing precision. Removing the regularisation term with $\lambda_{\text{reg}}=0$ results in a higher entanglement, which is expected since we do not penalise the network when changing other features other than implicitly through the pixel and feature loss. Training without the feature correlation relaxation in the regularising term ($\lambda_{\text{cor}}=0$) significantly increases the edit error as it constrains the model more. At the same time, we can also see that this model achieves a better entanglement score compared to SeFFeC, which is expected since this term allows for entanglement of correlated features in SeFFeC. Finally, we compare the results with SeFFeC 3x, which iteratively applies SeFFeC three times by feeding $w^{+}_{\text{edit}}$ back as the input to the transformer. As the base SeFFeC might over- or undershoot the desired target value for a semantic face feature, iteratively applying it multiple times can significantly reduce the edit error. However, this comes at the cost of an error increase for the other metrics.
Figure 5: Correlation matrix showing the Pearson correlation between two semantic face features in each cell. Correlations for all our features can be found in the supplementary material.
@@ 4.5 Correlation of the Face Features
The ablation results in Table 3 show that the relaxation of the regularising term based on feature correlations as defined in Equation 3 is important to increase edit precision by purposefully entangling correlated features. Figure 5 shows the Pearson correlation coefficient matrix for a subset of our features (see full correlation matrix with evaluation in supplementary material). We estimate these correlations by calculating all of our semantic face features for each face in the FFHQ dataset. We can see from the matrix that most features have almost no correlation, while only a small set of pairs are strongly correlated. For example, eye width and eyebrow width are strongly correlated, which makes sense intuitively as it is uncommon for humans to have small eyes with very wide eyebrows (and vice versa). We can see a similarly strong correlation between mouth width and nose width. Therefore, by reducing the loss penalty for the network when editing both features together, we give it more freedom in actually performing the change while also ensuring that the edited faces remain aesthetically plausible. We can also see that jaw width is correlated with every other feature related to width. As an increase in jaw width generally relates to a wider face, the other facial features can be scaled accordingly without receiving a high penalty. Similarly, a very wide face with small facial features is not very common. Another example for a strong correlation is between nose length and mouth height, which can be explained by the fact that the distance between the nose and the upper lip, called Philtrum, is typically around 2 cm with only slight variations [61]. Therefore, those two features are strongly correlated and change together to allow the Philtrum to stay within the natural range.
While most feature correlations can intuitively be explained by natural correlations in human faces, some correlations are likely due to a limitation of our absolute distance semantic face features. One such example is the correlation between chin length and mouth height. As both of these features are absolute distance measures, they are both sensitive to translations and rotations in y-direction. For example, if a head is rotated backwards, both of these features will decrease because chin and mouth move up in the image space, resulting in a false positive correlation. This suggest that when using more invariant features in future work, results can be improved even further.

### v3

@@ 4.1 Implementation Details
Input to our Transformer encoder is the sequence of vectors $w^{+}\in\mathbb{R}^{512\times 18}$ concatenated with the $512$ dimensional semantic face feature embedding extracted by the embedding layer. The encoder consists of four Transformer encoder layers with four self-attention heads [51] each. The output of the Transformer is input to a linear layer that outputs the semantic manipulation vector $s_{e}$. The scaling network takes the current and target face feature values, $m_{j}$ and $m_{j}^{t}$ for a feature $j$, and the feature’s embedding $e_{j}$ as input. It consists of three linear layers with 32, 32, and one hidden unit(s), respectively, with a ReLU activation function after the first two linear layers and outputs the scaling value $k$.
To train UP-FacE, we generate data batches by sampling random vectors $z\sim\mathcal{N}(0,1)$ from a standard normal distribution and generate the corresponding vectors $w^{+}$ and images $I$ using the mapping and synthesis network from a pre-trained StyleGAN2 [23] model. For each image, we then sample a face feature index $j$ we want to modify from a uniform distribution $j\sim\mathcal{U}(0,23)$ as well as a desired target feature value $m_{j}^{t}$ from a standard normal distribution $m_{j}^{t}\sim\mathcal{N}(0,1)$. Furthermore, we use the pre-trained SPIGA landmark detector [37] to extract landmarks and calculate the current face feature values $m_{j}$ for each generated image $I$. Since we sample $m_{j}^{t}$ from $\mathcal{N}(0,1)$ we normalise $m_{j}$ to follow a standard normal distribution as well:
@@
where $\mu$ and $\sigma$ are the mean and standard deviation for all face features $j$ calculated on the FFHQ dataset, which our StyleGAN2 model was pre-trained on. This ensures that our defined semantic face features are normalised to the same value range when calculating the loss. Using these generated data samples, we can run UP-FacE to predict $w^{+}_{\text{edit}}$, use the StyleGAN2 synthesis network to generate $I_{\text{edit}}$ and the SPIGA landmark detector to extract the new landmarks and calculate the predicted feature values $m_{j}^{p}$. We calculate the loss as defined in Equation 3 and propagate the gradients back through SPIGA and StyleGAN to finally update the parameters of UP-FacE. We optimise UP-FacE for $10^{5}$ steps using the Adam optimiser [25] with a learning rate of $2^{-5}$ and a batch size of 16. Through empirical testing, we set the weighting scalars of the loss function to $\lambda_{\text{pix}}=1$, $\lambda_{\text{feat}}=3$, $\lambda_{\text{SFF}}=0.005$, $\lambda_{\text{reg}}=0.1$ and $\lambda_{\text{cor}}=1$.
Figure 3: Example progressive edits performed with UP-FacE. UP-FacE allows to easily and deterministically perform high-quality progressive edits along many different semantic dimensions with explicit control over the desired target feature values. For each demonstration of the progressive edits, we also show the difference between the first and last image, highlighting which parts of the image changed.
Figure 4: Sample face editing results on a real face image of UP-FacE in comparison with the two state-of-the-art methods GANSpace [15] and InterFaceGAN [41]. The original image was inverted into latent space using the e4e framework [50], and subsequently, the degree of smiling was edited. Also shown are the difference images between the edited and original images. UP-FacE is the only method that allows for fine-grained and deterministic control of the degree of smiling without distortions.
@@ 4.2 Qualitative Results
Figure 1 shows sample sequential edits performed using UP-FacE. We first generated a face by sampling a random latent vector and decoded it with StyleGAN2 (left). Afterwards, we sequentially edited specific features by changing the current value to the desired target value, as indicated by the arrows. We can see that our method can perform multiple fine-grained edits on the desired features while preserving the visual appearance of the rest of the face, suggesting a high degree of disentanglement. In stark contrast to prior work, we can calculate the exact value of our face features at any point and set desired values to perform precise and user-predictable editing operations. This allows users to anticipate the outcome of an edit before its execution, e.g. setting the feature for mouth openness to a value of -1.1 always results in a closed mouth when normalised to a standard normal distribution. Note that it is possible to define reasonable bounds for each face feature and subsequently normalise the values to an arbitrary range as we did for our intuitive slider user interface. In this case, the smallest value is zero, indicating a closed mouth, and the largest value is one, indicating a fully opened mouth.
@@
Given that our features are on a continuous scale, UP-FacE naturally enables progressive edits as shown in Figure 3: UP-FacE can perform the desired changes with fine-grained control over the desired feature values. For each example, the figure also shows the difference between the first and the last image in the edit sequence. We can see that UP-FacE operates highly localised, as significant changes are only visible in the relevant regions. Other changes visible in the difference images are mostly due to high-frequency features, such as the outline of the face. Minor variations along these lines are visible in the difference images but cannot be noticed in the actual images.
@@
Overall, the examples in Figure 1 and Figure 3 show that our method can perform localised and precise face editing for many facial features. While the semantic face features defined in Table 1 only relate a small number of landmarks to each other, we can model complex face dynamics accurately. For example, while the definition for mouth openness only relates the y-coordinate of two landmarks from the centre of the mouth, the progressive editing results for this feature in Figure 3 show how the whole mouth opens naturally. We hypothesise that the main reason for the effectiveness of our semantic face features is within the implicit biases and correlations learned by the StyleGAN2 model. To move the two landmarks used for mouth openness apart, UP-FacE learns to open the whole mouth as this is the only possibility given the learned human face distribution of StyleGAN2.
@@
Figure 4 shows editing results on a real face that we inverted into StyleGAN2’s latent space using the e4e encoder [50]. We edited the facial expression twice: once to show less and once to show more smiling. As the results for UP-FacE in the middle row show, it is also possible to manipulate real faces with high precision, which is generally harder as the inverted latent code of the face might not lay within the well-defined regions of StyleGAN2. Since smiling is one of the attributes that previous methods can also control, we compare our results with GANSpace [15] and InterFaceGAN [41]. We can see that both methods fail to accurately reduce the degree of smiling and start to distort other parts of the face, which is most noticeable in the difference images. When increasing the degree of smiling, all three methods can accurately edit the face, while UP-FacE still operates more locally. Additionally, UP-FacE provides more fine-grained control over the mouth as it can edit the mouth shape and mouth openness separately, allowing us to add a smile without necessarily opening the mouth. Important to note is that UP-FacE did not require any manually labelled training data, while methods like InterFaceGAN rely on large face datasets with smiling attribute annotations.
Model Determ. Editing No. of Labels Accuracy FID
GANSpace 0 76.1% 36.89
InterFaceGAN 30.000 83.2% 32.58
UP-FacE 0 81.8% 27.15
@@ 4.3 Quantitative Results
We quantitatively evaluate UP-FacE on the smile edit benchmark [29, 28] and compare its performance against two strong baselines, GANSpace [15] and InterFaceGAN [41], using the same StyleGAN2 model for a fair comparison. The task of this benchmark is to convert faces with neutral expressions into smiling faces. The performance is measured with three metrics: attribute accuracy, which measures whether a face is smiling after editing using an attribute classifier pre-trained on the CelebA [30] dataset, Fréchet Inception Distance (FID) [18], calculated between 4000 edited images and the FFHQ dataset to evaluate image quality [23], and Identity Score (ID Score), which measures if the faces’ identity is preserved after editing by calculating the cosine similarity between embeddings extracted from a pre-trained ArcFace model [9]. To make a neutral face smile with UP-FacE  we adjust mouth shape and mouth openness as shown in Figure 4. The results of the smile edit benchmark are reported in Table 2. We find that UP-FacE significantly outperforms GANSpace and InterFaceGAN regarding FID and ID scores. While InterFaceGAN achieves a slightly higher attribute accuracy than UP-FacE, it is important to note that it was trained on the same 30.000 attribute labels as the classifier used to calculate the accuracy metric. In contrast, UP-FacE did not require any attribute labels and did not explicitly learn to manipulate the smiling attribute while still being competitive with InterFaceGAN and outperforming GANSpace. Furthermore, unlike previous methods, UP-FacE allows for precise and predictable control over face shape features. This could potentially be one reason for the significant improvement in FID and ID scores, as UP-FacE can automatically adjust the editing intensity based on the measured values of the face features. This allows UP-FacE  for example, to only slightly modify faces that are almost smiling, whereas previous methods would move a fixed amount along a semantic editing direction, which can lead to unnecessary over-editing and distortion.
Model Pixel     Error $\downarrow$
LIPIPS $\downarrow$
Edit     Error $\downarrow$ Entangle- ment $\downarrow$
$\lambda_{\text{pix}}=0$ 0.192 0.258 0.664 0.624
$\lambda_{\text{feat}}=0$ 0.041 0.045 0.551 0.238
$\lambda_{\text{reg}}=0$ 0.043 0.047 0.549 0.267
$\lambda_{\text{cor}}=0$ 0.040 0.041 0.575 0.225
UP-FacE 0.041 0.045 0.529 0.234
UP-FacE 3x 0.065 0.087 0.318 0.340
@@ 4.4 Ablation Experiments
We perform ablation experiments for different versions of UP-FacE and report four different metrics: pixel error, LIPIPS, edit error and entanglement. The pixel error is defined as the mean absolute error (MAE) between the original and the modified image. While we want this to be as small as possible, it should never be zero as this would indicate no change in the face. LIPIPS [62] is a perceptual similarity metric based on neural network features where a smaller value indicates higher similarity. The edit error is defined as the MAE between the predicted feature value $m_{j}^{\text{p}}$ and the target $m_{j}^{\text{t}}$ for the features $m_{j}$ that we actually want to change. Similarly, entanglement measures the MAE between $m_{j}^{\text{p}}$ and $m_{j}^{\text{t}}$ for all features that were not explicitly changed and should stay close to the original value. We report the results in Table 3, which we calculated by generating 10,000 random images and then performed edits on five random semantic features for each face.
We can see that the model is unstable during training without the pixel loss ($\lambda_{\text{pix}}=0$), which is reflected in the lower performance. Training without the feature loss ($\lambda_{\text{feat}}=0$) slightly increases the edit error, indicating that the feature supervision adds some flexibility compared to pure pixel supervision, improving editing precision. Removing the regularisation term with $\lambda_{\text{reg}}=0$ results in a higher entanglement, which is expected since we do not penalise the network when changing features other than implicitly through the pixel and feature loss. Training without the feature correlation relaxation in the regularising term ($\lambda_{\text{cor}}=0$) significantly increases the edit error as it constrains the model more. At the same time, we can also see that this model achieves a better entanglement score compared to UP-FacE, which is expected since this term allows for the entanglement of correlated features in UP-FacE. Finally, we compare the results with UP-FacE 3x, which iteratively applies UP-FacE three times by feeding $w^{+}_{\text{edit}}$ back as the input to the transformer. UP-FacE may sometimes over- or undershoot the desired target value for a semantic face feature. However, by iteratively applying the method multiple times, the edit error can be significantly reduced, resulting in more precise and user-predictable editing.
Figure 5: Correlation matrix showing the Pearson correlation between two semantic face features in each cell.
@@ 4.5 Correlation of the Face Features
The ablation results in Table 3 show that the relaxation of the regularising term based on feature correlations as defined in Equation 3 is important to increase edit precision by purposefully entangling correlated features. Figure 5 shows the Pearson correlation coefficient matrix for a subset of our features. We estimate these correlations by calculating all of our semantic face features for each face in the FFHQ dataset. We can see from the matrix that most features have almost no correlation, while only a small set of pairs are strongly correlated. For example, eye width and eyebrow width are strongly correlated, which makes sense intuitively as it is uncommon for humans to have small eyes with very wide eyebrows (and vice versa). We can see a similarly strong correlation between mouth width and nose width. Therefore, by reducing the loss penalty for the network when editing both features together, we give it more freedom to perform the change while ensuring that the edited faces remain aesthetically plausible. We can also see that jaw width is correlated with every other feature related to width. As an increase in jaw width generally relates to a wider face, the other facial features can be scaled accordingly without receiving a high penalty. Similarly, a wide face with small facial features is uncommon. Another example of a strong correlation is between nose length and mouth height, which can be explained by the fact that the distance between the nose and the upper lip, called Philtrum, is typically around 2 cm with only slight variations [60]. Therefore, those two features are strongly correlated and change together, allowing the Philtrum to stay within the natural range.
While most feature correlations can intuitively be explained by natural correlations in human faces, some correlations are likely due to a limitation of our absolute distance semantic face features. One such example is the correlation between chin length and mouth height. Both of these features are absolute distance measures, so they are both sensitive to translations and rotations in y-direction. For example, if a head is rotated backwards, both of these features will decrease because the chin and mouth move up in the image space, resulting in a false positive correlation. This suggests that results can be improved even further when using more invariant features in future work.